# Distribution Transformers: Fast Approximate Bayesian Inference With On-The-Fly Prior Adaptation

### v1

@@ Abstract
Machine Learning, ICML, Approximate Inference, Transformers
@@ 1 Introduction
Modern machine learning has transformed many data-rich domains (LeCun et al., 2015), yet critical applications like healthcare diagnostics and autonomous vehicle safety demand more than just predictive accuracy (Hüllermeier & Waegeman, 2019). These high-stakes settings, often characterized by limited or noisy data, require robust uncertainty estimation and principled integration of domain knowledge (Gawlikowski et al., 2021)—capabilities naturally addressed through probabilistic reasoning.
Bayesian inference is a cornerstone of probabilistic machine learning, encoding prior beliefs and providing the tools to incorporate new observations while providing principled uncertainty quantification; essential for critical decision-making under uncertainty. However, Bayesian inference is typically intractable, limiting its real-world adoption. To counteract this problem, the field of approximate inference has emerged, providing a range of methods: Variational inference (VI) methods (Blei et al., 2016) can approximate posterior densities but require expensive optimization procedures at inference time (as opposed to training time, relevant for later methods). While Markov Chain Monte Carlo (MCMC) (Christian & Casella, 2011) methods provide asymptotically-exact posterior samples, they neither directly estimate the density nor offer the rapid inference often required in practical applications.
Recent advances in deep learning have opened new avenues for Bayesian inference. Prior Fitted Networks (PFNs) (Muller et al., 2021) demonstrated that transformer (Vaswani et al., 2017) architectures can perform Bayesian inference through a single forward pass, dramatically reducing inference time compared to traditional methods. However, PFNs are limited by their dependence on a fixed prior, requiring costly retraining for each new prior distribution encountered. In practice, it is often desirable, and in some applications essential, to be able to change the prior rapidly at inference time.
While these methods have their strengths, there remains a clear need for approximate inference methods that can provide accurate posterior density estimates with minimal computational overhead while maintaining the flexibility to modify priors at inference time. To address this need, we introduce Distribution Transformers (DTs), a novel transformer-based architecture that combines the flexibility of VI with the efficiency of PFNs. Our method generates posterior density estimates through a single forward pass while handling a range of prior distributions from a given parametric family without retraining. Unlike traditional VI approaches that often rely on simple parametric families, DTs leverage a expressive distributional family that can universally approximate smooth densities. Figure 2 illustrates the architectural design of DTs. Fundamentally, DTs approximate the prior with a Gaussian Mixture Model (GMM) represented as an embedded sequence in the latent space of the transformer, then condition this GMM on observations using a permutation-equivariant transformer decoder, yielding a GMM approximation to the posterior.
DTs are trained using entirely sample-based supervised learning by defining a joint distribution over the prior, parametrized by $\phi$, the latent variable $x$, and the observations $z$. We demonstrate that this training approach yields an approximator of exact posterior prediction. This combination of speed, flexibility, and expressiveness enables accurate Bayesian inference in scenarios where existing methods face fundamental limitations. Concretely, DTs address the key challenges discussed above, providing accurate posterior density estimates for a wide range of priors. Like PFNs, our approach only requires a way to sample from the chosen prior family and an arbitrary likelihood model, and is widely applicable without significant task-specific analysis. We make the following contributions:
  1. 1.
    We introduce Distribution Transformers (DTs)—a novel architecture that leverages the power of universally-approximating Gaussian Mixture Models and Transformers to express arbitrary distribution-to-distribution mappings.
  2. 2.
    We show that DTs can be trained to perform more accurate approximate Bayesian inference, and up to orders of magnitude faster, than existing approximate inference methods, while also allowing on-the-fly updates to the prior at inference time without retraining, a never before accomplished feat.
  3. 3.
    We empirically demonstrate near-exact inference capabilities on challenging problems, including joint PPD-hyperposterior approximations for Gaussian Processes (GPs), parameter inference for quantum systems and a real-world sequential inference task, while remaining fast enough for real-time applications.
@@ 1.1 Related Work
[…]
Variational inference (VI) aims to approximate the posterior density through fitting a density of a particular family (typically a Gaussian) to the posterior by maximizing a variational lower bound on the evidence. Notable modern extensions include stochastic variational inference (SVI) (Hoffman et al., 2013), which utilizes backpropagation to optimize a stochastic approximation of the variational objective, black-box variational inference (Ranganath et al., 2014), which provides widely-applicable, model-agnostic tools for optimization of this variational objective, and amortized variational inference (AVI) (Kingma & Welling, 2014; Ganguly et al., 2023), which efficiently solves many variational inference problems in parallel through the learning of an inference network mapping from observations to approximate posteriors. These methods often provide acceptable estimates for the posterior density, but necessarily involve an expensive optimization routine at inference time, are not asymptotically exact, and often underestimate the posterior variance. Nonetheless, VI remains the most popular approximate inference technique in use today, and we will use a variant as a baseline in our empirical studies.
Recent neural approaches to approximate inference leverage the expressivity of deep learning architectures while aiming to maintain theoretical guarantees. Prior Fitted Networks (PFNs) (Muller et al., 2021) learn to perform approximate inference through sample-based supervised learning without the need for posterior densities (which are instead accessed via samples from the joint distribution over $x$ and $z$), enabling inference in a single forward pass. They provide rapid inference for changes in dataset, provide a posterior density estimate, and allow for almost arbitrary priors, but require complete retraining upon changes to the prior distribution, and use a Riemann Distribution, which can struggle to model smooth, heavy-tailed posteriors. As a close competitor to our method, we will also adopt a PFN baseline for our posterior studies.
Continuous normalizing flows (Rezende & Mohamed, 2015) learn a mapping from samples of some simple distribution to samples of the target posterior density, leveraging automatic differentiation to provide a flexible, closed-form density using a change of measure. However, the mapping they learn is fixed, meaning the prior is set at training time, and observations can only be incorporated via sufficient statistics of the simple distribution, limiting their suitability for our problem setting. Moreover, they must be constructed using only fully reversible components, which can limit expressivity. Classical approaches include the Laplace approximation (Tierney & Kadane, 1986) and expectation propagation (EP) (Minka, 2001). The Laplace approximation fits a Gaussian distribution based on the local geometry at the posterior mode, but lacks invariance to reparametrization. EP iteratively refines an isotropic Gaussian approximation to the posterior, but struggles with numerical stability.
Sampling-based methods, most notably Markov chain Monte Carlo (MCMC) methods, provide samples from the posterior which can be used to compute stochastic estimates for expectations with respect to said posterior. These are primarily based on access to the unnormalized posterior via the numerator of Bayes’ theorem. When a differentiable prior and likelihood model are available, the state-of-the-art No U-Turn Sampler (NUTS) (Hoffman & Gelman, 2011) can be used. NUTS is based on Hamiltonian Monte Carlo (HMC) (Brooks et al., 2012), which provides more efficient generation of posterior samples without problem-specific tuning of the algorithm. MCMC methods are asymptotically exact and almost universally applicable, but are typically expensive due to random walk behavior. They also do not provide direct access to a posterior density, and empirical density estimates can be sample inefficient.
Bayesian quadrature-based methods (Osborne et al., 2012) assign a probabilistic model, typically a GP, to the integrand of the evidence $p(z)$ (and in some cases the predictive posterior distribution PPD $p(y|x,z)=\int p(y|x)p(x\mathop{|}z)$), and use probabilistic numerical integration to estimate the evidence and PPD directly. These methods provide accurate uncertainty measures, and can handle noisy, black-box likelihood models, but are typically computationally expensive and scale poorly to high-dimensional problems, although recent advances in sparse-GPs (Binois & Wycoff, 2021) provide promising avenues to alleviate these scaling issues.
1.1.1 Sequential Inference
Sequential inference over time series requires dedicated methods, referred to as Bayesian filtering (when causality is required) or smoothing (when not). The extended Kalman filter (EKF) linearizes dynamics and observation models, then approximates the distribution over the latent state as a closed-form Gaussian. With modern automatic differentiation, this method is widely applicable and computationally efficient, but struggles in practice due to the overly restrictive assumptions induced by linearization and Gaussian distributions. Particle filters (PFs), or sequential Monte Carlo methods (Wills & Schön, 2023), maintain a population of particles which are distributed according to the posterior distribution each time step, but can be expensive and only provide an empirical density. The unscented Kalman filter (Julier & Uhlmann, 1997) maintains a minimal population of particles so is computationally efficient, but faces issues of numerical stability and also encodes a Gaussian assumption thus cannot handle multimodality. We compare our method against an EKF and a PF baseline.
@@ 2.1 Transformers
Two theoretical properties of transformers are central to our work. First, transformers are universal sequence-to-sequence approximators (Yun et al., 2019), capable of representing arbitrary mappings between sequences. Second, in the absence of positional encodings, transformers are permutation equivariant with respect to the input sequence—a property we will exploit in Section 3.
Our method specifically employs the transformer decoder architecture. This architecture extends the base transformer by incorporating global cross-attention layers, allowing each position in the input sequence to attend to a separate context sequence. This cross-attention mechanism provides a natural framework for conditioning sequence transformations on observed data.
@@ 2.2 Gaussian Mixture Models
Gaussian Mixture Models (GMMs) are flexible probability distributions whose density is a weighted sum of Gaussian components, i.e. $q(x)=\sum_{i}w_{i}\mathcal{N}(x;\mu_{i},\Sigma_{i})$ where $\mathcal{N}(x;\mu,\Sigma)$ is a Gaussian density over $x,w_{i}\in[0,1]$, $\sum_{i}w_{i}=1$, $\boldsymbol{\mu}_{i}\in\mathbb{R}^{n}$, and $\boldsymbol{\Sigma}_{i}\in\mathbb{S}_{++}^{n}$, illustrated in Figure 1. While GMMs are widely used in clustering (Dempster et al., 1977), and latent variable modelling (Bishop, 2006), we focus on their role as universal approximators of smooth probability distributions (Goodfellow et al., 2016; Calcaterra & Boldt, 2008)—a property we exploit in Section 3. This universality extends to distributions on compact domains under appropriate change of measure, also illustrated in Figure 1. While the idea of using GMMs for function approximation is not new to deep learning (for example Bishop (1994) proposes the use of a GMM to model uncertainty in the output of a neural network), the idea of operating end-to-end on a distribution represented as a GMM is novel.
@@
A key property of GMMs is their natural representation as an unordered sequence of component parameters. Specifically, a $k$-component GMM over $\mathbb{R}^{n}$ is parametrized by $\boldsymbol{\theta}=\{(w_{i},\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\}_{i=1}^{k}$. This representation is permutation invariant—the ordering of components does not affect the resulting distribution.
Fitting a GMM to a given probability distribution is non-trivial, with methods such as expectation maximization and variational methods suffering from similar problems to their approximate inference counterparts. We will show that DTs naturally provide a mapping from the parameters of a given distribution to an approximating GMM without introduction of additional model parameters.
@@ 2.3 Amortization
Amortized variational inference (AVI) (Ganguly et al., 2023; Kingma & Welling, 2014) exemplifies this paradigm, where a neural network is trained to approximate posterior distributions across multiple datasets, bypassing the need for iterative optimization for each new dataset. Similarly, PFNs (Muller et al., 2021) implicitly amortize inference through the learning of a mapping from observations to approximate posterior distributions without iterative updates. These examples highlight how amortization enables scalability and efficiency in otherwise computationally intensive workflows. As we shall show in Section 3, our work is also well-set in the context of amortized methods.
@@ 3 Distribution Transformers
     Estimate loss ^$\hat{\ell}^{\prime}_{\theta}=-\sum_{i=1}^{b}\log q_{\theta}(f(x_{i})\mathop{|}\phi_{i})+\log q_{\theta}(f(x_{i})\mathop{|}z_{i},\phi_{i})$;
@@
The proposed loss $l_{\theta}$ equals the expected KL-Divergence $\mathbb{E}_{p(\phi,z)}\left[\text{KL}\left[p(\cdot\mathop{|}z,\phi)\mathop{||}q_{\theta}(\cdot\mathop{|}z,\phi)\right]\right]$ between $p(\cdot\mathop{|}z,\phi)$ and $q_{\theta}(\cdot\mathop{|}z,\phi)$ up to an additive constant. Proof in Appendix A.
The loss $l_{\theta}$ can be calculated stochastically using samples from the joint distribution $p(\phi,x,z)$, alleviating any need to directly access or sample from the posterior density.
Finally, DTs can jointly approximate priors and posteriors without additional model parameters. Applying the unembedding to the prior sequence yields a GMM approximation $q_{\theta}(x\mathop{|}\phi)$, extending DTs to mappings $\Phi\times\mathcal{Z}\rightarrow\Theta\times\Theta$. To ensure consistency, we introduce a prior loss:
@@
leading to the combined objective: $\ell_{\theta}^{\prime}=\ell_{\theta}^{\text{prior}}+\ell_{\theta}$
@@ 4 Empirical Studies
We study the behaviour of our method in three settings: approximation of a tractable posterior, approximation of intractable posteriors, and a real-world sensor fusion problem posed as sequential inference. In the prior two settings, we benchmark our method against SVI (Hoffman et al., 2013), implemented in PyTorch (Paszke et al., 2019) with GPU parallelization, and PFNs (Muller et al., 2021), also implemented in PyTorch, fitting a Riemann distribution with the same number of model outputs as our DT. In the absence of a fixed prior, we train PFNs using the same sampling scheme as our method, effectively marginalizing out the meta-prior, leaving a less informative prior. We expect PFNs to perform well when the meta-prior is narrow and poorly when wide. For the latter experiment, we benchmark against the industry standard extended Kalman filter (EKF), and a particle filter (PF), again with GPU implementation. For the PF, we obtain a density via Gaussian kernel density estimation.

### v2

@@ 1 Introduction
Bayesian inference provides a principled route to uncertainty quantification and the incorporation of prior knowledge. In practice, however, repeatedly solving inference problems (e.g., across datasets, hyperparameters, or environments) is computationally expensive. Amortized Bayesian inference (ABI) addresses this by learning a mapping from observations to posterior approximations, yielding fast test-time inference. Recent transformer-based ABI methods have shown impressive single-pass inference for small-data regimes (Muller et al., 2021; Hollmann et al., 2022, 2025).
Yet two limitations persist. First, most ABI models fix the prior during training; changing the prior at test time typically requires retraining or fine-tuning. Second, methods that offer some prior flexibility usually do not preserve family structure between prior and posterior, hindering sequential composition (filtering/smoothing), where the posterior must become the next-step prior. Complementary lines in amortized SBI (Cranmer et al., 2020) and sensitivity-aware prior amortization (Elsemüller et al., 2023) underscore the need for flexible priors, but do not maintain the at-least approximate conjugacy needed for recursive updates.
We introduce Distribution Transformers (DTs), a transformer-based architecture that (i) performs single-pass amortized inference, (ii) supports prior amortization across a family of priors at test time, and (iii) approximates conjugacy between prior and posterior, enabling clean sequential composition when needed. Concretely, DTs embed priors as a tokenized sequence (representing a Gaussian Mixture Model, or GMM, approximation), condition on observations with a permutation-equivariant decoder, and output a posterior in the same family. Training spans a distribution over priors, enabling prior-amortization.
Concretely, we make the following contributions:
  • A unified framework for prior-flexible, approximately conjugate amortized inference, bridging the gap between ABI and sequential inference.
  • A practical transformer architecture that tokenizes distributions and performs posterior updates within the same parametric family.
  • Empirical results on both static benchmarks (matching or exceeding PFN/TabPFN- and VI-style baselines) and sequential tasks (where preserving conjugacy is critical), demonstrating accuracy and speed.
@@ 1.1 Related Work
Amortized Bayesian Inference.
Amortization has emerged as a key paradigm to accelerate Bayesian inference across repeated tasks by transferring expensive optimization costs to an offline training phase. Classical amortized variational inference (AVI) approaches (Kingma & Welling, 2013; Ganguly et al., 2023b) learn inference networks that map directly from observations to approximate posteriors, but typically rely on restrictive variational families and do not support flexible prior specification. Recent extensions have explored amortization in simulation-based inference (SBI), leveraging normalizing flows (Rezende & Mohamed, 2015; Papamakarios et al., 2017), neural ratio estimation (Greenberg et al., 2019; Miller et al., 2022), and hybrid MCMC-amortized approaches (Salimans et al., 2015; Gabrié et al., 2022), but these remain tied to fixed priors at training time. Prior-amortized methods have only recently been proposed, for example prior-amortized neural posterior estimation for reflectometry inversion (Starostin et al., 2025), and sensitivity-aware amortized inference (Elsemüller et al., 2023), highlighting both the importance and nascent state of this line of work. Relatedly, BayesFlow provides an amortized SBI framework based on invertible neural networks and tooling for practical workflows (Radev et al., 2020).
Transformers for Bayesian Inference.
Transformers have recently shown promise as amortized inference engines. Prior-Fitted Networks (PFNs) (Muller et al., 2021) demonstrated that transformers can approximate posterior distributions in a single forward pass, amortizing inference over datasets. However, PFNs assume a fixed prior, and their Riemannian output distribution struggles with smooth or heavy-tailed posteriors. Follow-up work extended PFNs to tabular data (TabPFN) (Hollmann et al., 2022), scaling to larger contexts and small-data regimes with improved accuracy (Hollmann et al., 2025), and to time-series forecasting (Hoo et al., 2024). Nonetheless, these methods remain restricted to fixed priors [^1] A rudimentary form of prior amortization can be carried out by providing prior parameters as additional observations, but this is restrictive and a side-effect not a directly-intended feature.. In parallel, amortized in-context Bayesian inference methods (Mittal et al., 2025; Reuter et al., 2025) investigate whether transformers can learn posterior inference directly from prompts, but still lack mechanisms for prior adaptation.
Meta-Learning and Neural Processes.
Our work connects to neural processes (Garnelo et al., 2018a, b), which condition on context sets to predict function values, and their transformer-based extensions (Kim et al., 2019; Nguyen & Grover, 2022). While these frameworks amortize inference across tasks, they typically model predictive distributions over data rather than explicit posteriors over latent variables. ACE (Chang et al., 2024) generalized this direction to incorporate latents and flexible priors, and is thus complementary to our work. Other amortized meta-learning approaches (Wu et al., 2020; Iakovleva et al., 2020) have proposed shared amortized inference networks across tasks, but again without explicit prior adaptation.
Simulation-Based Inference.
SBI methods (Cranmer et al., 2020; Lueckmann et al., 2017; Papamakarios et al., 2019) provide amortized posterior approximations for complex simulators, with recent advances leveraging transformers and diffusion models (Gloeckler et al., 2024; Sharrock et al., 2022; Wildberger et al., 2023). These works are highly expressive and achieve state-of-the-art inference in scientific applications, yet typically assume a fixed prior distribution or limited parametric families. The lack of efficient prior amortization limits their applicability in scenarios requiring frequent prior updates, such as sensitivity analysis or robust sequential decision-making.
Sequential Inference.
Classical Bayesian filtering methods–such as the Kalman filter (Kalman, 1960), unscented Kalman filter (Julier & Uhlmann, 1997), and particle filters (Doucet et al., 2001; Wills & Schön, 2023)–remain the dominant approaches to sequential inference. While computationally efficient, they are either constrained to Gaussian assumptions or expensive particle-based representations, and provide no general amortization across tasks or priors. Despite the centrality of sequential inference in real-world applications, modern amortized inference frameworks (PFNs, TabPFN, ACE) have largely neglected this dimension.
Positioning.
In summary, prior work has established the effectiveness of amortization in approximate Bayesian inference, and the suitability of transformers for fast, in-context posterior approximation. However, flexible prior amortization, particularly in the context of sequential inference, remains largely unaddressed. Our work introduces Distribution Transformers as a principled solution, combining (i) the expressive universality of Gaussian mixtures, (ii) transformer-based amortization across priors, and (iii) applicability to sequential Bayesian filtering—providing capabilities unmatched by existing PFN, TabPFN, ACE, or simulation-based inference approaches.
@@ 2.1 Transformers
Two theoretical properties of transformers are central to our work. First, transformers are universal sequence-to-sequence approximators (Yun et al., 2019), capable of learning arbitrary mappings between sequences. Second, in the absence of positional encodings, transformers are permutation equivariant with respect to the input sequence—a property we exploit in Section 3.
Our method specifically employs the transformer decoder architecture. This architecture extends the base transformer by incorporating global cross-attention layers, allowing each token in the input sequence to attend to a separate context sequence. This cross-attention mechanism provides a natural framework for conditioning sequence transformations on observed data.
@@ 2.2 Gaussian Mixture Models
Gaussian Mixture Models (GMMs) are flexible probability distributions whose density is a weighted sum of Gaussian components, i.e. $q(x)=\sum_{i}w_{i}\mathcal{N}(x;\mu_{i},\Sigma_{i})$ where $\mathcal{N}(x;\mu,\Sigma)$ is a Gaussian density over $x,w_{i}\in[0,1]$, $\sum_{i}w_{i}=1$, $\boldsymbol{\mu}_{i}\in\mathbb{R}^{n}$, and $\boldsymbol{\Sigma}_{i}\in\mathbb{S}_{++}^{n}$, illustrated in Figure 1. While GMMs are widely used in clustering (Dempster et al., 1977), and latent variable modelling (Bishop, 2006), we focus on their role as universal approximators of smooth probability distributions (Goodfellow et al., 2016; Calcaterra & Boldt, 2008)—a property we exploit in Section 3. This universality extends to distributions on compact domains under appropriate change of measure, also illustrated in Figure 1. While the idea of using GMMs for function approximation is not new to deep learning (for example (Bishop, 1994) proposes the use of a GMM to model uncertainty in the output of a neural network), the idea of operating end-to-end on a distribution represented as a GMM is novel.
@@
A key property of GMMs is their natural representation as an unordered sequence of component parameters. Specifically, a $k$-component GMM over $\mathbb{R}^{n}$ is parametrized by $\boldsymbol{\theta}=\{(w_{i},\boldsymbol{\mu}_{i},\boldsymbol{\Sigma}_{i})\}_{i=1}^{k}$. This representation is permutation invariant—the ordering of components does not affect the resulting distribution. Fitting a GMM to a given probability distribution is non-trivial, with methods such as expectation maximization and variational methods suffering from similar problems to their approximate inference counterparts. We will show that DTs naturally provide a mapping from the parameters of a given distribution to an approximating GMM without introduction of additional model parameters.
@@ 2.3 Amortization
Amortized variational inference (AVI) (Ganguly et al., 2023a; Kingma & Welling, 2014) exemplifies this paradigm, where a neural network is trained to approximate posterior distributions across multiple datasets, bypassing the need for iterative optimization for each new dataset. Similarly, PFNs (Muller et al., 2021) implicitly amortize inference through the learning of a mapping from observations to approximate posterior distributions without iterative updates. These examples highlight how amortization enables scalability and efficiency in otherwise computationally intensive workflows. As we shall show in Section 3, our work is also well-set in the context of amortized methods.
@@ 3 Distribution Transformers
  Estimate loss $\hat{\ell}^{\prime}_{\theta}=-\sum_{i=1}^{b}\log q_{\theta}(f(x_{i})\mathop{|}\phi_{i})+\log q_{\theta}(f(x_{i})\mathop{|}z_{i},\phi_{i})$;
@@
The proposed loss $l_{\theta}$ is equal to the expected KL-Divergence $\mathbb{E}_{p(\phi,z)}\left[\text{KL}\left[p\mathop{||}q_{\theta}\right]\right]$ between $p(\cdot\mathop{|}z,\phi)$ and $q_{\theta}(\cdot\mathop{|}z,\phi)$ up to an additive constant. Proof in Appendix A.
The loss $l_{\theta}$ can be estimated using samples from the joint distribution $p(\phi,x,z)$, alleviating any need to directly access or sample from the posterior density.
Finally, DTs can jointly approximate priors and posteriors without additional model parameters. Applying the unembedding to the prior sequence yields a GMM approximation $q_{\theta}(x\mathop{|}\phi)$, extending DTs to mappings $\Phi\times\mathcal{Z}\rightarrow\Theta\times\Theta$. To ensure consistency and a shared latent space pre- and post-conditioning, we introduce a prior loss:
@@
leading to the combined objective: $\ell_{\theta}^{\prime}=\ell_{\theta}^{\text{prior}}+\ell_{\theta}$ .
@@ 4 Empirical Studies
We study the behaviour of our method in three settings: approximation of a tractable posterior, approximation of intractable posteriors, and a real-world sensor fusion problem posed as sequential inference. In the prior two settings, we benchmark our method against SVI (Hoffman et al., 2013), implemented in PyTorch (Paszke et al., 2019) with GPU parallelization, and PFNs (Muller et al., 2021), also implemented in PyTorch, fitting a Riemann distribution with the same number of model outputs as our DT. We share out code through an anonymised link  [^2] https://anonymous.4open.science/r/distribution-transformers-450D/. In the absence of a fixed prior, we train PFNs using the same sampling scheme as our method, effectively marginalizing out the meta-prior, leaving a less informative prior. We expect PFNs to perform well when the meta-prior is narrow and poorly when wide, as the effective prior used by the PFN is the marginalisation of the prior family with respect to the meta prior and so is close to the true prior only in the former case. For the latter experiment, we benchmark against the industry standard extended Kalman filter (EKF), and a particle filter (PF), again with GPU implementation. For the PF, we obtain a density via Gaussian kernel density estimation.